how fast are you going if you travel 45 miles in 50 minutes?
Answer by user_dude2008(1861)

50 min = 50/60 = 5/6 hrs


d=rt
45=r(5/6)
45*6=5r
270=5r
270/5=r
54=r

r=54 mph